OpenStudy (anonymous):

John’s piggy bank contained 7 dimes and 3 quarters. He pulled out 1 coin without looking. Without replacing the first coin, John then pulled out a second coin. What is the probability that both coins were dimes? Express your answer in simplest form.

OpenStudy (anonymous):

\[\frac{ ? }{ ? }\]

OpenStudy (anonymous):

@inowalst

OpenStudy (anonymous):

@JohnKnows

OpenStudy (anonymous):

@mavagirl552000

OpenStudy (anonymous):

@CaseyBabe

OpenStudy (anonymous):

plz help

OpenStudy (perl):

the first dime, how many ways can that occur,

OpenStudy (anonymous):

7 dimes total or 10 coins total

OpenStudy (perl):

ok, so the chance to pick a dime will be 7/10, right?

OpenStudy (perl):

now once you pick a dime, there are 6 dimes left (out of 9 coins). multiply probabilities

OpenStudy (perl):

$$ \Large \frac{7}{10}\cdot \frac{6}{9} $$

OpenStudy (anonymous):

oh ok so it is\[\frac{ 42 }{ 90 }\]

OpenStudy (anonymous):

or\[\frac{ 21 }{ 45 }\]